OpenStudy (anonymous):

solve a/2=a-6/3a-9-1/3

OpenStudy (zzr0ck3r):

use more parenthesis \(\frac{a}{2}=\frac{a-6}{3a-9}-\frac{1}{3}\\?\)

OpenStudy (anonymous):

your supposed to get two answers which is 1 and 2

OpenStudy (zzr0ck3r):

do I have it correct?

OpenStudy (anonymous):

yeah you do

OpenStudy (zzr0ck3r):

it is very hard to see what you are asking without the parentheses.

OpenStudy (zzr0ck3r):

ok

OpenStudy (anonymous):

i need help

OpenStudy (zzr0ck3r):

\(\frac{a}{2}=\frac{a-6}{3a-9}-\frac{1}{3}=\frac{a}{2}=\frac{a-6}{3(a-3)}-\frac{1}{3}\) so our \(LCM=2*3*(a-3)\) so multiply everything by our \(LCM\) and we get \(2*3*(a-3)\frac{a}{2}=2*3*(a-3)\frac{a-6}{3(a-3)}-2*3*(a-3)\frac{1}{3}=\\\cancel{2}*3*(a-3)\frac{a}{\cancel{2}}=2*\cancel{3*(a-3)}\frac{a-6}{\cancel{3(a-3)}}-2*\cancel{3}*(a-3)\frac{1}{\cancel{3}}=\\3(a-3)a=2(a-6)-2(a-3)\). you with me?

OpenStudy (anonymous):

yeah

OpenStudy (zzr0ck3r):

ok \(3(a-3)a=2(a-6)-2(a-3)\iff3a^2-9a=2a-12-2a+6 \iff \\3a^2-9a=-6\iff3a^2-9a+6=0\)

OpenStudy (zzr0ck3r):

can you solve that?

OpenStudy (anonymous):

im not sure I followed that right

OpenStudy (anonymous):

could you start from the beginning

OpenStudy (anonymous):

to say im confused on the way u set up the problem